OpenStudy (tiffany_rhodes):

A homebuyer wishes to take out a mortgage of $500,000 for a 30 year period. What monthly payment is required if the interest rate is 10% ?

ganeshie8 (ganeshie8):

use the attached formula

ganeshie8 (ganeshie8):

okay, it wont let me attach :/ here it is : http://prntscr.com/38lur4

OpenStudy (tiffany_rhodes):

It's alright! Thanks! Is the periodic interest rate 10% in my case? I'm a little hazy with loan/mortgage terminology. :P

OpenStudy (mathmale):

Cool, ganeshie8! I once used this formula often, but have long since forgotten it. However, either you or I could find it again by looking up "amortization formula."

ganeshie8 (ganeshie8):

good q :) how many months does one year has ?

OpenStudy (tiffany_rhodes):

12 :D

ganeshie8 (ganeshie8):

10% is the annual interest rate...

ganeshie8 (ganeshie8):

when they dont specify, always take it as annual

ganeshie8 (ganeshie8):

yes, so periodic interest rate, i = r/n i = 0.1 / 12

OpenStudy (tiffany_rhodes):

alright. So would my formula look like this: \[500,000 * 0.0083(1+0.0083)^(30) /(1+0.0083)^(30) -1\]

OpenStudy (tiffany_rhodes):

Where 500,000 is the principal; 0.0083 is .1/12 as the periodic interest rate; and 30 years is the "length of the loan in years" ?

ganeshie8 (ganeshie8):

ur interpretation is right, but the numbers u have plugged are wrong

ganeshie8 (ganeshie8):

P = 500000 t = 30 years r = 0.1 n = 12 i = r/n = 0.1/12

ganeshie8 (ganeshie8):

you forgot to multiply 30*12 in the exponent...

OpenStudy (tiffany_rhodes):

Oh okay, I was confused because it said "length of the loan in YEARS" so I assumed 30

OpenStudy (tiffany_rhodes):

thank you!

ganeshie8 (ganeshie8):

t = length of loan in years = 30 n = number of compounding intervals in a year = 12 (monthly)

ganeshie8 (ganeshie8):

since every month you're paying back some money, the loan needs to adjust monthly accordingly

ganeshie8 (ganeshie8):

^^use wolfram for these lengthy calculations

OpenStudy (tiffany_rhodes):

Alright, thank you for your help @ganeshie8 !

ganeshie8 (ganeshie8):

np :)